OpenStudy (anonymous):

URGENTLY needed AOB is a diameter of the circle and AC = BC, then angleCAB is equal to

OpenStudy (anonymous):

Friends i have done , but doubt ki mera solution is correct or not Since, AOB is the diameter, ∠ACB = 90 (angles in a semi-circle is a right angle) ΔABC is an isosceles triangle [AB = BC(Given)] ∴ ∠CAB = ∠ABC (isosceles triangle property) but, ∠CAB + ∠ABC + ∠ACB = 180 (angle sum property of a triangle) ∠CAB + ∠ABC + 90= 180 [∠ACB = 90(proved)] 2∠CAB + 90 = 180 [∠CAB = ∠ABC (proved)] 2∠CAB = 180 - 90 = 90 ∠CAB = 90/2 = 45 ∠CAB = 45

OpenStudy (anonymous):

yes ur solution is correct

OpenStudy (anonymous):

@Ahmad_Shadab thank you so much

OpenStudy (anonymous):

 This illustration also indicates that your answer is correct.